OpenStudy (anonymous):

What is the simplified form of (2/x^2 + x) - (1/x) ?

OpenStudy (anonymous):

\[\frac{ 2}{x^2 + x}-\frac{1}{x}\]

OpenStudy (anonymous):

\[x^2+x=x(x+1)\] hint

OpenStudy (anonymous):

how a common denominator

OpenStudy (anonymous):

Why would you assume it is equal to zero?